OpenStudy (anonymous):

jack has a hamburger stand setup at a school fete. the ingredients to make the hamburgers cost $55. if she sells each hamburger for $2, how many does she need to sell to make a profit of $65?

OpenStudy (anonymous):

60

OpenStudy (anonymous):

55+65=120 120/2=60

OpenStudy (anonymous):

solve using algebra

OpenStudy (anonymous):

(55+65)/2=x 120/2=x 60=x

OpenStudy (anonymous):

thanx alot